# Thermodynamics of Modified Chaplygin Jacobi And Modified Chaplygin Abel Gas: Stability Analysis and Observational Constraints

### v1

@@ 5   Conclusion
Our investigation presents a comprehensive analysis of the thermodynamic characteristics exhibited by two cosmological models: the Modified Chaplygin Jacobi Gas (MCJG) and Modified Chaplygin Abel Gas (MCAG) in the context of universal evolution. Through our analysis, we have derived a novel thermal equation of state expressing pressure as a function of temperature and volume, $P=P(T,V)$.
The research demonstrates that both gas models exhibit distinct behavioral patterns at different cosmic scales. At minimal volumes, they manifest as pressureless systems, while at expanded volumes, they maintain a consistent negative pressure. This behavior effectively illustrates the cosmic transition from a matter-dominated era to one governed by cosmological constant dynamics. Our thermodynamic framework establishes specific constraints within which the thermal equation of state operates, particularly noting its temperature dependence. The thermal range is bounded by $0<T<T^{*}$, where $T^{*}$ represents the maximum attainable temperature for both gas models at minimal volumes. The investigation of cosmic acceleration reveals significant insights through various cosmological parameters. The deceleration parameter $q$ exhibits positive values at small volumes but transitions to negative values as volume increases, confirming a shift from decelerated to accelerated expansion in the late universe. The dark energy equation of state parameter $\omega$ maintains negative values with an increasing trend during cosmic expansion for both models. The density parameter $\Omega$ corroborates this evolutionary pattern, indicating a transition from matter to dark energy dominance near $z\approx 0.5$.
@@
These findings collectively establish that both the Modified Chaplygin Jacobi and Modified Chaplygin Abel gas models demonstrate thermodynamic stability while maintaining classical stability against perturbations, providing a robust framework for understanding cosmic evolution. We have also performed observational data anaysis using CC+BAO and Pantheon+SH0ES datasets. The information criterion shown in Tab. 2 indicates that both the model are observationally viable. In summary, both the MCJG and MCAG models exhibit similar trends in parameter values between the Pantheon+SH0ES and CC+BAO datasets, but with notable differences in the magnitude of certain parameters, particularly $\Omega_{m0}$, $\omega_{m}$, $\alpha$, and $V_{0}$. These variations reflect differing assumptions about the roles of matter, dark energy, and potential interactions in the two models.

### v2

@@ 5   Conclusion
Our investigation presents a comprehensive analysis of the thermodynamic characteristics exhibited by two cosmological models: the Modified Chaplygin-Jacobi Gas and Modified Chaplygin-Abel Gas in the context of universal evolution. Through our analysis, we have derived a novel thermal equation of state expressing pressure as a function of temperature and volume, $P=P(T,V)$. The research demonstrates that both gas models exhibit distinct behavioral patterns at different cosmic scales. At minimal volumes, they manifest as pressure-less systems, while at expanded volumes, they maintain a consistent negative pressure. This behavior effectively illustrates the cosmic transition from a matter-dominated era to one governed by cosmological constant dynamics. Our thermodynamic framework establishes specific constraints within which the thermal equation of state operates, particularly noting its temperature dependence. The thermal range is bounded by $0<T<T^{*}$, where $T^{*}$ represents the maximum attainable temperature for both gas models at minimal volumes. The investigation of cosmic acceleration reveals significant insights through various cosmological parameters. The deceleration parameter $q$ exhibits positive values at small volumes but transitions to negative values as volume increases, confirming a shift from decelerated to accelerated expansion in the late universe. The dark energy equation of state parameter $\omega$ maintains negative values with an increasing trend during cosmic expansion for both models. The density parameter $\Omega$ corroborates this evolutionary pattern, indicating a transition from matter to dark energy dominance near $z\approx 0.5$.
@@
These findings collectively establish that both the Modified Chaplygin-Jacobi and Modified Chaplygin-Abel gas models demonstrate thermodynamic stability while maintaining classical stability against perturbations, providing a robust framework for understanding cosmic evolution. We have also performed observational data analysis using CC+BAO and Pantheon+SH0ES datasets. The information criterion shown in Tab. 2 indicates that both the model are observationally viable. In summary, both the MCJG and MCAG models exhibit similar trends in parameter values between the Pantheon+SH0ES and CC+BAO datasets, but with notable differences in the magnitude of certain parameters, particularly $\Omega_{m0}$, $\omega_{m}$, $\alpha$, and $V_{0}$. These variations reflect differing assumptions about the roles of matter, dark energy, and potential interactions in the two models.